# Approximation schemes for covering and packing mixed-integer programs with a fixed number of constraints

### v1

@@ 6 Conclusion and further research
This paper presents approximation schemes for various mixed-integer problems with a fixed number of constraints. By analyzing the structure of the extremes points of the associated polytope of such problems, we decompose it into multidimensional knapsack cover instances with a single continuous variable per dimension. By leveraging approximation algorithms from the knapsack literature, we obtain a PTAS for p. Utilizing a similar methodology, an FPTAS and an $\epsilon$-approximate formulation is presented for the one-dimensional case. For the one-dimensional case with uniform bounds, we propose a perfect compact formulation. To obtain these results, some classical knapsack results are extended to incorporate continuous variables. These results are especially valuable as they provide a framework to design algorithms and tight formulations for similarly structured problems such as fixed-charge, facility location, and supplier selection problems. We also derive analogous results for the packing and assignment variants of p.
@@ Appendix A Appendix
where $\bar{f}$ and $\bar{v}$ are both positive, $0<\bar{w}_{i}\leq\bar{d}$ for all $i\in[\eta]$, and $0<\bar{c}\leq\bar{d}$. Let $OPT^{*}$ be the value of the LP relaxation of 1-mkc. Assume, without loss of generality, that $\bar{f}_{1}\geq\bar{f}_{2}\geq\ldots\geq\bar{f}_{\eta}$. In this appendix, we prove the following theorem using an approach and structure that adheres closely to what is shown by bienstockTighteningSimpleMixedinteger2012.

### v2

@@ 7 Conclusion and further research
This paper presents approximation schemes for various mixed-integer problems with a fixed number of constraints. By analyzing the structure of the extremes points of the associated polytope of such problems, we decompose it into multidimensional knapsack cover instances with a single continuous variable per dimension. By leveraging approximation algorithms from the knapsack literature, we obtain a PTAS for p. Utilizing a similar methodology, an FPTAS and an $\epsilon$-approximate formulation is presented for the one-dimensional case. For the one-dimensional case with uniform bounds, we propose a perfect compact formulation. To obtain these results, some classical knapsack results are extended to incorporate continuous variables. These results are especially valuable as they provide a framework to design algorithms and tight formulations for similarly structured problems. We also derive analogous results for the packing and more general variants of p.
@@ Appendix A Appendix
where $\bar{f}$ and $\bar{v}$ are both positive, $0<\bar{w}_{i}\leq\bar{d}$ for all $i\in[\eta]$, and $0<\bar{c}\leq\bar{d}$. Let $OPT^{*}$ be the value of the LP relaxation of 1-mkc. Assume, without loss of generality, that $\bar{f}_{1}\geq\bar{f}_{2}\geq\ldots\geq\bar{f}_{\eta}$. In this appendix, we prove the following theorem using an approach and structure that adheres closely to what is shown by Bienstock and McClosky (2012).